OpenStudy (anonymous):

solve explain x/2 = 2/3 (x + 3/8)

OpenStudy (marigirl):

we are working with fractions here.. \[\frac{ x }{ 2 }=\frac{ 2 }{ 3 }(x+\frac{ 3 }{ 8 })\]

OpenStudy (marigirl):

first lets expand the brackets on the right side \[\frac{ x }{ 2 }=\frac{ 2x }{ 3 }+\frac{ 6 }{ 24 }\] are you okay with the step above?

OpenStudy (marigirl):

then what I did was collect like terms, \[\frac{ x }{ 2 }-\frac{ 2x }{ 3 }=\frac{ 6 }{ 24 }\] then minus the fractions on the left \[\frac{ 3x }{ 6 }-\frac{ 4x }{ 6 }=\frac{ 6 }{ 24 }\] \[\frac{ -x }{ 6 }=\frac{ 6 }{ 24 }\] multiply both sides by 6\[-x=\frac{ -36 }{ 24 }\] simplify \[x=\frac{ -3 }{ 2 }\]